OpenStudy (anonymous):

NEED HELP THANKS :) After how many months will Rita pay one hundred dollars in simple annual interest on a credit card that charges twelve percent on her five hundred dollars balance? 2 months 4 months fourteen months twenty months

OpenStudy (anonymous):

) I = Prt 100 = (500)(0.12)t 100 = 60t 100/60 = t t = 1 2/3 years = 20 months (fourth choice)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

your welcome ;-)